Two cards are drawn without replacement from an ordinary deck of 52 playing cards. What are the odds in favor of drawing two queens
Answer by Alan3354(69443)

Two cards are drawn without replacement from an ordinary deck of 52 playing cards. What are the odds in favor of drawing two queens
------------------
There are 4 queens at the start, so the odds are 4/52 = 1/13.
The odds of the 2nd Q are 3/52.
1/13 x 3/52 = 3/676